A mother is 21 years older than her son. 6 years
from now exactly, the mother will be exactly 5 times
as old as the child. Where's the child's dad?

Answer by Edwin McCravy(20056)

A mother is 21 years older than her son.
Let the son be S years old
Then the mother is S+21

6 years from now exactly, the mother will be exactly 5 times
as old as the child.
S+21+6 = 5(S+6)
  S+27 = 5S+30
    -3 = 4S
  -3/4 = S  

So the son is negative three-fourths of a year old, which means
he has not been born yet.  Since three fourths of a year
before he is born is 9 months before he is born, he is just
being conceived.

Where's the child's dad?

Having sex with his mother. LOL
Edwin